The length of a rectangle is 10 more than five times the width. If the perimeter is 152, find the length and the width
Answer by waynest(281)

length = 5w + 10
p = 152
p = 2w + 2(5w + 10)
152 = 2w + 10w + 20
152 - 20 = 12w + 20 - 20
132 = 12w
132/12 = 12w/12
11 = w
check:
152 = 2(11) + 2(5(11) + 10)
152 = 22 + 2(55 + 10)
152 = 22 + 2(65)
152 = 22 + 130
152 = 152